### v3

# SONAR: A Synthetic AI-Audio Detection Framework and Benchmark

Abstract
Recent advances in Text-to-Speech (TTS) and Voice-Conversion (VC) using generative Artificial Intelligence (AI) technology have made it possible to generate high-quality and realistic human-like audio. This introduces significant challenges to distinguishing AI-synthesized speech from the authentic human voice and could raise potential issues of misuse for malicious purposes such as impersonation and fraud, spreading misinformation, deepfakes, and scams. However, existing detection techniques for AI-synthesized audio have not kept pace and often exhibit poor generalization across diverse datasets. In this paper, we introduce SONAR, a synthetic AI-Audio Detection Framework and Benchmark, aiming to provide a comprehensive evaluation for distinguishing cutting-edge AI-synthesized auditory content. SONAR includes a novel evaluation dataset sourced from 9 diverse audio synthesis platforms, including leading TTS providers and state-of-the-art TTS models. It is the first framework to uniformly benchmark AI-audio detection across both traditional and foundation model-based deepfake detection systems. Through extensive experiments, we reveal the generalization limitations of existing detection methods and demonstrate that foundation models exhibit stronger generalization capabilities, which can be attributed to their model size and the scale and quality of pretraining data. Additionally, we explore the effectiveness and efficiency of few-shot fine-tuning in improving generalization, highlighting its potential for tailored applications, such as personalized detection systems for specific entities or individuals. Code and dataset are available at https://github.com/Jessegator/SONAR
1 Introduction
Recent advances in Text-to-Speech (TTS) and Voice-Conversion (VC) using Artificial Intelligence (AI) technology have made it possible to generate high-quality and realistic human-like audio efficiently [1, 2, 3, 4]. This introduces significant challenges in distinguishing AI-synthesized speech from the authentic human voice and could raise potential misuse for malicious purposes such as impersonation and fraud, spreading misinformation, and scams. For example, a deep fake AI voice of the US President Joe Biden was recently utilized in robocalls to advise them against voting [^1] https://www.cnn.com/2024/01/22/politics/fake-joe-biden-robocall/index.html, demonstrating how deepfakes can significantly manipulate public opinions and influence presidential elections. In response to such risks, the US Federal Communications Commission (FCC) now deems robot calls for election as illegal, which underscores the urgent need for enhanced detection of AI-synthesized audio.
While TTS models are advancing rapidly, AI-synthesized audio detection techniques are not keeping pace. First, previous studies [5, 6] have highlighted the lack of generalization and robustness in these detection methods. Second, existing detection models [7, 6, 8, 9, 10] often take advantage of different audio features and evaluation datasets, complicating the comparison of their detection effectiveness. Third, a comprehensive evaluation to determine the effectiveness of these detection methods against the latest TTS models has not been conducted. This gap in research leaves a significant challenge in developing reliable detection techniques that can effectively counter the growing sophistication of AI-generated audio.
To address the aforementioned research gap and explore the strengths and limitations of existing AI-synthesized audio detection methods, especially those with increasingly advanced TTS models, we present a synthetic AI-Audio Detection Framework and Benchmark, coined as SONAR. This framework aims to provide a comprehensive evaluation for distinguishing state-of-the-art AI-synthesized auditory content. Our study benchmarks the state-of-the-art fake audio detection models using a newly collected fake audio dataset that includes a variety of synthetic speech audios sourced from diverse cutting-edge TTS providers and TTS models. We further investigate the potential of enhancing the generalization capabilities of these detection models from different perspectives. The main contributions of our work can be summarized as follows.
  • We introduce a novel evaluation dataset specifically designed for audio deepfake detection. This dataset is sourced from 9 diverse audio synthesis platforms, including those from leading TTS service providers and state-of-the-art TTS models. To the best of our knowledge, this dataset is by far the largest collection of fake audio generated by the latest TTS models.
  • SONAR is the first comprehensive framework to benchmark AI-audio detection uniformly across advanced TTS models. It covers 5 state-of-the-art traditional and 6 foundation-model-based audio deepfake detection models.
  • Leveraging SONAR, we conduct extensive experiments to analyze the generalizability limitations of current detection methods. Our findings reveal that foundation models demonstrate stronger generalization capabilities than traditional models. We further explore factors that may contribute to this improved generalization, such as model size and the scale and quality of pre-training data.
  • We further explore the potential of few-shot fine-tuning to enhance the generalization of detection models. Our empirical results demonstrate the effectiveness and efficiency of this approach, highlighting its potential for tailored applications, such as personalized detection systems for specific entities or individuals.
2 Related work
Text-to-Speech synthesis. Human voice synthesis is a significant challenge in the field of AI. State-of-the-art TTS synthesis approaches such as VALLE [4], AudioBox [1], VoiceBox [11], NaturalSpeech3 [4], and YourTTS [3] have demonstrated the possibility of generating high-quality, human-realistic audio with generative models trained on large datasets. Current TTS models can be classified into two primary categories: cascaded and end-to-end methods. Cascaded TTS models [12, 13, 14] typically employ a pipeline involving an acoustic model and a vocoder utilizing mel spectrograms as intermediary representations. To address the limitations associated with vocoders, end-to-end TTS models [15, 16] have been developed to jointly optimize both the acoustic model and vocoder. In practical applications, it is preferable to customize TTS systems to generate speech in any voice with limited accessible data. Consequently, there is increasing interest in zero-shot multi-speaker TTS techniques [17, 3, 2].
AI-synthesized audio detection. Recent advancements in AI technology have significantly enhanced the ability to generate high-quality and realistic audio, calling for an urgent need for more robust and reliable detection methods. Several datasets have been developed to support research in this area. The ASVspoof challenges [18, 19, 20] are among the most notable, offering comprehensive datasets that cover a variety of attack vectors, including replay attacks, voice conversion, and directly synthesized audio. These resources aim to facilitate thorough evaluations of countermeasures against various spoofing techniques. In addition, newer datasets such as WaveFake [21] and LibriSeVoc [22] provide fake audio samples generated with state-of-the-art vocoders, offering diverse distributions to enhance the development of deepfake audio detection systems. By comparison, the In-the-Wild dataset [6] targets real-world applications by collecting deepfake audios from publicly accessible sources, capturing the complexity and diversity of manipulations encountered in everyday environments. Similarly, the SingFake dataset [6] focuses on the detection of synthetic singing voices, presenting unique challenges due to the musical content and variation in vocal expressions. These datasets are crucial for developing and testing next-generation AI-synthesized audio detection systems, pushing the boundaries of what is achievable in identifying and mitigating the threats posed by advanced audio synthesis technologies.
Building upon these datasets, a significant body of research has focused on distinguishing AI-generated audio from genuine audio by designing advanced model architectures [8, 10, 7, 9] tailored for extracting different levels of representations of speech data for audio deepfake detection. Additionally, recent works [23, 24, 25] have leveraged speech foundation models for audio deepfake detection tasks. For instance, [23] and [24] fine-tune Wav2Vec2 [26] models on the ASVspoof dataset, while [25] uses Whisper as a front-end to extract audio features and trains various detection models based on these features, achieving state-of-the-art detection performance on the corresponding test datasets. However, none of these models have been evaluated on audio generated by the latest text-to-speech (TTS) models, leaving a gap in understanding their effectiveness against the most recent advancements in synthetic audio generation.
3 Evaluation dataset generalization and collection
Figure 1: Overview of SONAR. Left: Audio deepfake data generation and collection. Right: Benchmark evaluation.
Leveraging a set of diverse and high-quality speech data synthesis APIs and models, we create an evaluation dataset for synthetic AI-audio detection. Our approach incorporates two strategies: data generation and data collection. Our dataset includes AI-generated speech and audio from nine distinct sources. We perform speech data generation using one cutting-edge TTS service provider, OpenAI, and two open-sourced APIs, xTTS [27] and AudioGen [28]. For speech data collection, we leverage six state-of-the-art TTS models including Seed-TTS [29], VALL-E [4], PromptTTS2 [30], NaturalSpeech3 [31], VoiceBox [11], FlashSpeech [2]. Table 1 presents the details of our dataset generated by different audio generation models. We next detail our methods of generating and collecting these datasets.
Table 1: Overview of our dataset with fake audios generated by various models. AudioGen lacks speaker and language information as it focuses on environmental sounds. The trainset sizes for OpenAI and Seed-TTS are unavailable due to the use of proprietary data. * denotes the samples that are directly collected from their demo page or provided test set due to the unavailability of their model checkpoints.
Model Samples Avg duration (s) Avg. pitch (Hz) Std. pitch (Hz) Languages Trainset size(H) Year
PromptTTS2* 25 9.86 126.49 46.27 English 44K 2023
NaturalSpeech3* 32 5.25 143.86 53.94 English 60K 2024
VALL-E* 95 4.86 133.41 56.54 English 60K 2023
VoiceBox* 104 10.28 114.09 37.89 English, German, French, 60K 2023
Portuguese, Polish, Spanish
FlashSpeech* 118 7.57 129.30 54.77 English 44.5K 2024
AudioGen 100 5.00 199.45 72.94 - 7K 2022
xTTS 600 5.67 164.67 95.20 English 2.7K 2023
Seed-TTS* 600 4.91 117.28 36.85 English, Mandarin - 2024
OpenAI 600 4.11 126.89 54.89 English - 2024
Data generation. Our dataset generation involves OpenAI, xTTS, and AudioGen. Specifically, OpenAI currently provides voice choices from 6 different speakers. Using ChatGPT,we generate 100 different text prompts of varying lengths for each speaker, resulting in a total of 600 synthetic speech audios. xTTS supports synthetic speech generation given text prompts and reference speech. We select 6 speakers from the LibriTTS dataset [32] as the reference speech and also generate 600 text prompts with ChatGPT for each speaker, resulting in 600 synthetic speech audios. AudioGen can generate the corresponding environmental sound given a textual description of the acoustic scene. With AudioGen, we use ChatGPT to generate 100 text descriptions of the environment and background and obtain 100 AI-synthesized environmental sounds. Figure 1 (left) illustrates the data generation and collection process.
Data collection. To evaluate the effectiveness of various detection systems against the state-of-the-art TTS models, we also collect fake speech audio from Seed-TTS, VALL-E, PromptTTS2, NaturalSpeech3, VoiceBox, and FlashSpeech. Seed-TTS provides a test dataset [^2] https://github.com/BytedanceSpeech/seed-tts-eval consisting of fake audio samples generated by it. Due to the unavailability of pre-trained weights of the other 5 models, we extract the synthesized speech data directly from their demo pages. Specifically, speech audios from VALL-E include variations in emotions and acoustic environment. PromptTTS2 presents fake audio samples with various attributes such as gender, speed, pitch, volume, and timbre. NaturalSpeech3 also includes fake audio samples generated with various attributes such as speeds and emotions and contains fake speech audio samples obtained with voice conversion. VoiceBox provides fake audio samples that feature cross-lingual and expressive audio styles. FlashSpeech includes a set of high-quality fake audios obtained both from speech generation and voice conversion.
To summarize, leveraging the details outlined above, we generate and collect a comprehensive evaluation dataset, encompassing a total of 2274 AI-synthesized audio samples produced by various TTS models. To the best of our knowledge, our dataset is by far the largest collection of fake audio generated by the latest TTS models. Note our motivation for collecting this dataset is for evaluation purposes. Additionally, we only include fake audio samples in this dataset since genuine audio samples can be easily collected from various sources (e.g., internet, self-recording, publicly available datasets, etc.). However, for convenience of evaluation, we also provide an equal number of real speech audio data sampled from the LibriTTS [32] clean-test set.
We believe the collected dataset can serve as a valuable asset for evaluating existing audio deepfake detection models.
4 Benchmarking AI-Audio Detection Models
In this section, we first detail the model, dataset, and evaluation metrics setup for benchmarking. Then, we present the results of evaluating detection models on existing audio deepfake datasets to assess their generalizability across datasets. We next benchmark their detection performance on our proposed dataset and provide analysis for potential model generalization improvement.
4.1 Benchmarking setup
Model architectures. SONAR incorporates 11 models, including 5 state-of-the-art traditional audio deepfake detection models featuring various levels of input feature abstraction and 6 foundation models. Specifically, for the former, SONAR includes (1) AASIST [7], which processes raw waveform directly and utilizes graph neural networks and incorporates spectro-temporal attention mechanisms. (2) RawGAT-ST [9], which employs spectral and temporal sub-graphs along with a graph pooling strategy. (3) RawNet2 [8], which is a hybrid model combining CNN and GRU.(4) Spectrogram(Spec.)+ResNet [6], which transforms the audio to linear spectrogram using a 512-point Fast Fourier Transform (FFT) with a hop size of 10 ms. The spectrogram is then inputted into ResNet18 [33]. (5) LFCC-LCNN [10], which converts audio into Linear-Frequency Cepstral Coefficients (LFCC) for input into a CNN model. Specifically, 60-dimensional LFCCs are extracted from each utterance frame, with frame length set to 20ms and hop size 10ms. It extracts speech embedding directly from raw audio. These models collectively cover a broad spectrum of feature types and architectures, facilitating a detailed examination of their performance in deepfake audio detection applications. For foundation models, SONAR includes (1) Wave2Vec2 [26], which is pre-trained on 53k hours of unlabeled speech data. (2) Wave2Vec2BERT [34], which is pre-trained on 4.5M hours of unlabeled speech data covering more than 143 languages. (3) HuBERT [35], which is pretrained-on 60k hours of speech data. (4) CLAP [36], who is trained on a variety of audio-text pairs. (5) Whisper-small [37], and (6) Whisper-large [37]. Both Whispers are pre-trained on 680K hours of speech data covering 96 languages.
Public datasets for training and testing. We consider three benchmark datasets for deepfake audio detection model training and testing as they are commonly used in the literature [38, 39]. Wavefake [21] collects deepfake audios from six vocoder architectures, including MelGAN [40], FullBand-MelGAN, MultiBand-MelGAN [41], HiFi-GAN [42], Parallel WaveGAN [43], and WaveGlow [44]. It consists of approximately 196 hours of generated audio files derived from the LJSPEECH [45] dataset. Similar to wavefake, LibriSeVoc [22] collects deepfake audios from six state-of-the-art neural vocoders including WaveNet[46], WaveRNN [47], Mel-GAN [41], Parallel WaveGAN [43], WaveGrad [48] and DiffWave [49] to generate speech samples derived from the widely used LibriTTS speech corpus [32], which is often utilized in text-to-speech research. Specifically, it consists of a total of 208.74 hours of synthesized samples. In-the-wild [5] comprises genuine and deepfake audio recordings of 58 politicians and other public figures gathered from publicly accessible sources, including social networks and video streaming platforms.
For LibriSeVoc, we follow the official train-validation-test splits, which are approximately 60%, 20%, and 20%, respectively. For Wavefake, we partition the data generated by each vocoder into training, validation, and testing subsets at ratios of 70%, 10%, and 20%, respectively. To address the class imbalance and mitigate potential evaluation bias, we further downsample LibriSeVoc and WaveFake test datasets, and In-the-Wild datasets, resulting in a balanced dataset with a real-to-fake ratio of 1:1.
Evaluation metrics. To provide a comprehensive evaluation of the detection performance of audio deepfake models, we adopt (1) Equal Error Rate (EER), which is defined as the point on the ROC curve, where the false positive rate (FPR) and false negative rate (FNR) are equal and is commonly used to assess the performance of binary classifications tasks, with lower values indicating better detection performance. (2) Accuracy evaluates the overall correctness of the detection model’s predictions and is defined as the ratio of correctly predicted data to the total data. To ensure consistency with the EER and provide more intuitive results, we set the threshold for accuracy at the EER point, meaning the accuracy reflects the model’s performance when the FPR equals the FNR. (3) AUROC (Area Under the Receiver Operating Characteristic) provides a measure of the model’s ability to distinguish between classes across different decision thresholds, providing a more comprehensive view of its discriminative power across varying conditions. An AUROC score of 1.0 indicates perfect classification, while a score of 0.5 indicates performance no better than random guessing.
Note that the test datasets are class-balanced, and the accuracy score is calculated using the EER threshold. Thus, we omit F1, precision, and recall scores from our evaluation results in the paper, though SONAR provides these metrics as well.
4.2 Results and analysis
4.2.1 How well can detection models generalize across datasets?
We first train all models on Wavefake training dataset and then evaluate the models on its own test set, LibriSeVoc test set, and In-the-wild dataset. Table 2 presents the evaluation results. Particularly, we make the following interesting observations.
Table 2: Generalization across existing audio deepfake datasets. All models are trained/finetuned on the Wavefake training set. Green and orange indicate the best and second-best performance, respectively.
Model Wavefake LibriSeVoc In-the-wild
Accuracy AUROC EER(%) Accuracy AUROC EER(%) Accuracy AUROC EER(%)
LFCC-LCNN 0.9984 0.9999 0.153 0.7429 0.8239 25.71 0.5 0.4786 99.2
Spec.+ResNet 0.9924 0.9924 0.076 0.7577 0.8495 24.233 0.4685 0.4723 53.148
RawNet2 0.9416 0.9592 5.839 0.5119 0.5332 48.807 0.5321 0.5393 46.792
RawGATST 0.9988 0.9999 0.115 0.8307 0.9203 16.925 0.6418 0.7015 35.816
AASIST 0.9992 0.9999 0.076 0.886 0.9534 11.397 0.7272 0.7975 27.277
CLAP 0.9996 0.9999 0.038 0.8296 0.9019 24.763 0.3013 0.2252 69.871
Whisper-small 0.9935 0.9997 0.649 0.9345 0.9837 6.551 0.821 0.9025 17.899
Whisper-large 0.9962 0.9992 0.381 0.9572 0.9901 4.279 0.8848 0.9552 11.518
Wave2Vec2 0.9874 0.9987 1.259 0.9705 0.9953 2.953 0.8733 0.9323 12.669
HuBERT 0.9931 0.9996 0.687 0.986 0.9991 1.401 0.9164 0.9653 8.362
Wave2Vec2BERT 0.9996 0.9999 0.038 0.9902 0.9991 0.984 0.9232 0.979 7.676
Speech foundation models exhibit stronger generalizability. As shown in Table  2, when evaluated on the test set of Wavefake, all models demonstrate near-perfect performance across the three metrics. This can be attributed to the similarity between the test set and the training data. However, when tested on the LibriSeVoc and In-the-wild datasets, models such as LFCC-LCNN, Spec.+ResNet, RawNet2, RawGATST, and AASIST struggle to generalize effectively. This performance gap indicates significant overfitting to the training data, despite these models being specifically designed for audio deepfake detection tasks. In contrast, speech foundation models consistently display stronger generalizability. Notably, Wave2Vec2BERT achieves the highest generalizability, which may be attributed to its large-scale and diverse pretraining data. Pretrained on 4.5 million hours of unlabeled audio in more than 143 languages, Wave2Vec2BERT benefits from both scale and diversity. This suggests that a well-designed self-supervised model trained on diverse speech data can extract general and discriminative features, making it more applicable across different datasets for audio deepfake detection. It is important to note that CLAP, unlike other speech foundation models, does not generalize well across datasets. This is likely due to its primary focus on environmental audio data during pretraining, resulting in the extraction of irrelevant features for speech audio. This observation underscores that not all foundation models are equally suited for audio deepfake detection tasks.
Generalizability may increase with model size. In Table 2, it can be observed that Whisper-large always outperforms Whisper-small across all three datasets. In particular, on the LibriSeVoc test set, Whisper-large achieves accuracy, AUROC, and EER of 0.9572, 0.9901, 4.279%, respectively, which improves by 2.27%, 0.64%, and 2.272%, than that of Whisper-small. This trend is more evident in the In-the-wild dataset, which is closer to real-world scenarios since this dataset consists of speech data sourced from the internet. Specifically, Whisper-large achieves accuracy, AUROC, and EER of 0.8848, 0.9552, and 11.518%, respectively, which improves by 6.381%, 5.27%, and 6.381%, than that of Whisper-small. Further investigation will be made in Section 4.2.3
4.2.2 Results on SONAR dataset
We further evaluate all detection models on the proposed dataset. Table 3(a), Table 3(b), and Table 3(c) present the accuracy, AUROC, and EER of different detection models on our proposed SONAR dataset as described in Sec 3.
Table 3: Evaluation on SONAR dataset. Green and orange indicate the best and second-best performance, respectively.
(a) Accuracy ($\uparrow$).
Model PromptTTS2 NaturalSpeech3 VALL-E VoiceBox FlashSpeech AudioGen xTTS Seed-TTS OpenAI Average
LFCC-LCNN 0.5200 0.7500 0.6211 0.8462 0.7034 0.4600 0.7433 0.3058 0.5000 0.6055
Spec.+ResNet 0.5600 0.5000 0.5684 0.5481 0.6356 0.6800 0.8450 0.4167 0.6783 0.6036
RawNet2 0.6800 0.3125 0.4211 0.5385 0.4915 0.2600 0.6533 0.3733 0.3500 0.4534
RawGATST 0.8000 0.5312 0.6842 0.8173 0.5424 0.2400 0.6567 0.5833 0.4900 0.5939
AASIST 0.8400 0.5312 0.7789 0.8750 0.6610 0.6900 0.7300 0.6567 0.5150 0.6975
CLAP 0.5600 0.4688 0.6421 0.5288 0.6017 0.2500 0.4800 0.4000 0.3233 0.4727
Whisper-small 0.8800 0.5625 0.7158 0.7404 0.5678 0.8000 0.8050 0.5983 0.1883 0.6509
Whisper-large 1.000 0.6562 0.7895 0.7885 0.7288 0.8400 0.9033 0.5933 0.2900 0.7322
Wave2Vec2 0.9600 0.6875 0.8210 0.9327 0.8136 0.9900 0.7333 0.8683 0.5175 0.8138
HuBERT 1.0000 0.7500 0.9158 0.9712 0.9407 1.0000 0.8767 0.8900 0.5658 0.8789
Wave2Vec2BERT 1.0000 0.9062 0.9474 0.9712 0.9237 0.9700 0.9867 0.6017 0.7833 0.8989
(b) AUROC ($\uparrow$).
Model PromptTTS2 NaturalSpeech3 VALL-E VoiceBox FlashSpeech AudioGen xTTS Seed-TTS OpenAI Average
LFCC-LCNN 0.5696 0.7666 0.6967 0.9106 0.7945 0.4559 0.8163 0.2452 0.0967 0.5947
Spec.+ResNet 0.6064 0.4941 0.6217 0.5858 0.6891 0.7293 0.9205 0.4003 0.7450 0.6436
RawNet2 0.6944 0.2422 0.3695 0.6210 0.5203 0.3030 0.7210 0.3120 0.2940 0.4530
RawGATST 0.8704 0.5439 0.7490 0.8989 0.5742 0.2050 0.7317 0.6065 0.4795 0.6288
AASIST 0.9248 0.6172 0.8479 0.9433 0.7485 0.7466 0.8265 0.6893 0.5259 0.7633
CLAP 0.5712 0.4434 0.7223 0.5155 0.6533 0.1777 0.5114 0.3544 0.2407 0.4655
Whisper-small 0.9776 0.5762 0.8050 0.8400 0.6446 0.8284 0.8915 0.6326 0.108 0.7004
Whisper-large 1.0000 0.6992 0.9063 0.8552 0.7933 0.8926 0.9690 0.6558 0.2327 0.7782
Wave2Vec2 0.9952 0.7515 0.8751 0.9674 0.8438 0.9987 0.7931 0.9205 0.4881 0.8482
HuBERT 1.0000 0.8174 0.9719 0.9953 0.9871 1.0000 0.9496 0.9531 0.5585 0.9148
Wave2Vec2BERT 1.0000 0.9658 0.9860 0.9906 0.9666 0.9826 0.9980 0.6165 0.8607 0.9290
(c) EER(%) ($\downarrow$).
Model PromptTTS2 NaturalSpeech3 VALL-E VoiceBox FlashSpeech AudioGen xTTS Seed-TTS OpenAI Average
LFCC-LCNN 48.000 25.000 37.895 15.385 29.661 54.000 25.667 69.5 99.333 44.938
Spec.+ResNet 44.000 50.000 43.158 45.192 36.441 32.000 15.500 58.333 32.167 39.643
RawNet2 32.000 68.750 57.895 46.154 50.848 74.000 34.667 62.667 65.000 54.665
RawGATST 20.000 46.875 31.580 18.269 45.763 76.000 34.330 41.667 51.000 40.609
AASIST 16.000 46.875 22.105 12.500 33.898 31.000 27.000 34.333 48.500 30.246
CLAP 44.000 53.125 35.789 47.115 39.831 75.000 52.000 60.000 67.667 52.725
Whisper-small 12.000 43.750 28.421 25.962 43.220 20.000 19.500 40.167 81.167 34.910
Whisper-large 0.000 34.375 21.053 21.154 27.119 16.000 9.667 40.667 71.000 26.782
Wave2Vec2 4.000 31.250 17.895 6.731 18.644 1.000 26.667 13.167 48.333 18.632
HuBERT 0.000 25.000 8.421 2.885 5.932 0.000 12.333 11.000 43.500 12.119
Wave2Vec2BERT 0.000 9.375 5.263 2.885 7.627 3.000 1.333 39.833 21.667 10.109
Speech foundation models can better generalize on the SONAR dataset, but still not good enough. As presented in Table 3(a), speech foundation models again exhibit better generalizability on the fake audio samples generated by the latest TTS models. For instance, AASIST achieves 0.6975 average accuracy across audios generated by cutting-edge TTS models, which is the best performance among the traditional detection models. In contrast, speech foundation models Whisper-large, Wave2Vec2, HuBERT, and Wave2Vec2BERT achieve an average accuracy of 0.7322, 0.788, 0.8789, and 0.8989, respectively, which is higher than AASIST by 3.47%, 9.05%, 18.14%, and 20.14%, respectively. More specifically, even though Wave2Vec2BERT and HuBERT are only fine-tuned on Wavefake dataset, for PromptTTS2, VALL-E, VoiceBox, FalshSpeech, AudioGen, and xTTS, Wave2Vec2BERT can reach accuracies of 1.0, 0.9062, 0.9474, 0.9712, 0.9237, 0.97, and 0.9867, respectively, and HuBERT can achieve 1.0, 0.9158, 0.9712, 0.9407, 1.0 0.8767, and 0.89, respectively, demonstrating their potential capability of extract more distinguishable features compared to other models. It is also worth noting that Wave2Vec2BERT achieves an accuracy of 0.9062 on NaturalSpeech3, while all other models can only reach that $\leq 0.75$.
It is still challenging for detection models to correctly classify synthesized audio samples, especially those generated by the most advanced TTS service providers. While Wave2Vec2BERT achieves an overall average accuracy of 0.8989, it only reaches 0.6017 on Seed-TTS and 0.7833 on OpenAI. A similar pattern is also evident with HuBERT, Wave2Vec2, Whisper-large, and Whisper-small, which achieve just 0.5658, 0.4342, 0.29, and 0.1883 accuracy on OpenAI, respectively. This performance disparity is likely due to OpenAI and Seed-TTS having more advanced model architectures and being trained on proprietary, self-collected data, leading to higher-quality and more realistic speech generation. We will explore potential strategies to enhance their detection performance in Section 4.2.4. Overall, these results not only indicate that no single model consistently outperforms across all datasets but also underscore the ongoing difficulty in detecting synthesized audio from cutting-edge TTS systems, especially those developed by the most advanced TTS service providers. This highlights a huge gap between the rapid evolution of TTS technologies and the effectiveness of current audio deepfake detection methods, emphasizing the urgent need for the development of more robust and reliable detection algorithms.
Additionally, it is noteworthy that, compared to speech foundation models, the accuracy of all five traditional detection models on the AudioGen dataset, which consists of synthesized environmental sounds, remains relatively low. Specifically, LFCC-LCNN, Spec.+ResNet, RawNet2, RawGATST, and AASIST achieve accuracies of 0.46, 0.68, 0.26, 0.24, and 0.69, respectively. In contrast, Whisper-small, Whisper-large, Wave2Vec2, HuBERT, and Wave2VecBERT attain significantly higher accuracies of 0.8, 0.84, 0.99, 1.0, and 0.97, respectively. This discrepancy may be due to traditional detection models being trained exclusively on speech data, which limits their generalization to audio from different distributions. In comparison, foundation models demonstrate greater robustness to out-of-distribution audio samples. An exception to this is CLAP, which is an audio foundation model pre-trained on a variety of environmental audio-text pairs and only achieves an accuracy of 0.25 on AudioGen. Similar to previous results, it’s possibly due to the fact that its full-weight fine-tuning on speech data may have compromised its ability to effectively recognize environmental sounds, resulting in poor performance.
4.2.3 Can generalizability increase with model size?
Building on the observation that Whisper-large consistently outperforms Whisper-small, we extend our analysis with controlled experiments on the entire Whisper model family. Specifically, the Whisper family comprises five different model sizes: Whisper-tiny, Whisper-base, Whisper-small, Whisper-medium, and Whisper-large. Table 4 presents the number of model parameters of them. Specifically, each model is fine-tuned on the Wavefake training dataset using the same hyperparameters. Our results show that as model size increases, the generalizability of the models improves as well.
Table 4: Whisper model sizes.
Model #Params
Whisper-tiny 39M
Whisper-base 74M
Whisper-small 244M
Whisper-medium 769M
Whisper-large 1550M
Table 5 presents the detection performance of the Whisper models across the Wavefake, LibriSeVoc, and In-the-wild datasets. First, Whisper-tiny, despite its smaller size, still outperforms or achieves comparable detection performance to traditional detection models (recall Table 2) on the LibriSeVoc test set. This again validates the finding that foundation models exhibit stronger generalizability for audio deepfake detection tasks, even in their smallest configurations.
Table 5: Generalization across existing audio deepfake datasets. All Whisper models are trained/finetuned on the Wavefake training set. Green and orange indicate the best and second-best performance, respectively.
Model Wavefake LibriSeVoc In-the-wild
Accuracy AUROC EER(%) Accuracy AUROC EER(%) Accuracy AUROC EER(%)
Whisper-tiny 0.9839 0.9985 1.603 0.8557 0.9307 14.426 0.498 0.5 50.203
Whisper-base 0.9908 0.9996 0.916 0.9163 0.9734 8.368 0.7398 0.8124 26.024
Whisper-small 0.9935 0.9997 0.649 0.9345 0.9837 6.551 0.821 0.9025 17.899
Whisper-medium 0.9962 0.9999 0.381 0.944 0.985 5.604 0.8572 0.9288 14.277
Whisper-large 0.9962 0.9992 0.381 0.9572 0.9901 4.279 0.8848 0.9552 11.518
Table 6: Evaluation on SONAR dataset. Green and orange indicate the best and second-best performance, respectively.
(a) Accuracy ($\uparrow$).
Model PromptTTS2 NaturalSpeech3 VALL-E VoiceBox FlashSpeech AudioGen xTTS Seed-TTS OpenAI Average
Whisper-tiny 0.8000 0.3438 0.6947 0.6442 0.4661 0.73 0.6517 0.5067 0.0833 0.5467
Whisper-base 0.8400 0.4375 0.6947 0.6731 0.6017 0.6800 0.6550 0.4800 0.1117 0.5749
Whisper-small 0.8800 0.5625 0.7158 0.7404 0.5678 0.8000 0.8050 0.5983 0.1883 0.6509
Whisper-medium 0.96 0.6250 0.7895 0.8077 0.7119 0.8000 0.8400 0.5517 0.2183 0.7005
Whisper-large 1.000 0.6562 0.7895 0.7885 0.7288 0.8400 0.9033 0.5933 0.2900 0.7322
(b) AUROC ($\uparrow$).
Model PromptTTS2 NaturalSpeech3 VALL-E VoiceBox FlashSpeech AudioGen xTTS Seed-TTS OpenAI Average
Whisper-tiny 0.9136 0.2998 0.7436 0.7144 0.4886 0.7660 0.7239 0.5033 0.0454 0.5776
Whisper-base 0.9296 0.4326 0.7512 0.7482 0.6548 0.7505 0.7167 0.5152 0.041 0.6155
Whisper-small 0.9776 0.5762 0.8050 0.8400 0.6446 0.8284 0.8915 0.6326 0.108 0.7004
Whisper-medium 0.9984 0.6279 0.886 0.8578 0.7950 0.8640 0.9215 0.5858 0.1567 0.7437
Whisper-large 1.0000 0.6992 0.9063 0.8552 0.7933 0.8926 0.969 0.6558 0.2327 0.7782
(c) EER(%) ($\downarrow$).
Model PromptTTS2 NaturalSpeech3 VALL-E VoiceBox FlashSpeech AudioGen xTTS Seed-TTS OpenAI Average
Whisper-tiny 20.000 65.625 30.526 35.577 53.390 27.000 34.833 49.333 91.667 45.328
Whisper-base 16.000 56.250 30.526 32.692 36.831 32.000 34.500 52.000 88.833 42.811
Whisper-small 12.000 43.750 28.421 25.962 43.220 20.000 19.500 40.667 81.167 34.965
Whisper-medium 4.000 37.500 21.053 19.231 28.814 20.000 16.000 44.833 78.167 29.955
Whisper-large 0.000 34.375 21.053 21.154 27.119 16.000 9.667 40.167 71.000 26.726
Second, as the model size increases from Whisper-tiny to Whisper-large, both accuracy and AUROC improve significantly across the LibriSeVoc and In-the-wild datasets. Whisper-large achieves an accuracy of 95.72% and an AUROC of 0.9901 on LibriSeVoc, surpassing Whisper-tiny by 10.07% in accuracy. A more evident pattern can be observed on the In-the-wild dataset, where Whisper-large outperforms Whisper-tiny by 38.48% in accuracy. Furthermore, the Equal Error Rate (EER) decreases as the model size increases, indicating that larger models are not only more accurate but also better at minimizing both false positives and false negatives.
We also evaluate the Whisper family on the proposed SONAR dataset. Table 6(a), Table 6(b), Table 6(c) present the corresponding Accuracy, AUROC, and EER(%). A similar trend can also be observed. In Table 6(a), the accuracy of the Whisper models shows a clear upward trend as the model size increases from Whisper-tiny to Whisper-large. Whisper-tiny achieves an average accuracy of 0.5467, while Whisper-large reaches the highest average accuracy of 0.7322. Notably, Whisper-large performs best on almost all datasets, particularly with TTS models such as PromptTTS2, NaturalSpeech3, VALL-E, and OpenAI, highlighting its better generalizability. Additionally, Whisper-large’s performance is higher on challenging datasets like Seed-TTS and OpenAI, which are known for their high-quality synthesis. The smaller models (e.g., Whisper-tiny and Whisper-base), on the other hand, struggle to generalize effectively, particularly on datasets such as OpenAI, where the accuracy drops to 0.0833 for Whisper-tiny.
The results highlight the scalability of the Whisper models: larger models demonstrate better generalization across diverse test sets, underscoring the importance of model capacity in tackling challenging out-of-distribution data, such as audio generated by advanced TTS models.
4.2.4 On the effectiveness and efficiency of few-shot fine-tuning to improve generalization
Despite the challenges in generalizing across different datasets, we investigate whether there exist efficient solutions that can enhance models’ detection performance on those challenging subsets from SONAR dataset. To this end, we conduct a case study on Wave2Vec2BERT and HuBERT, as these models perform relatively poorly on the OpenAI and SeedTTS datasets but demonstrate competitive performance on other subsets. Specifically, we generate 100 additional fake audio samples using the OpenAI TTS API and randomly select another 100 fake audio samples from the SeedTTS test set for few-shot fine-tuning. Our study yields several interesting findings.
Figures 2(a) and 2(b) present the results of fine-tuning Wave2Vec2BERT and HuBERT using varying numbers of samples from OpenAI. Before fine-tuning, Wave2Vec2BERT and HuBERT only achieve accuracies of 0.7833 and 0.5658, respectively. Notably, with only 10 shots of fake speech data, Wave2Vec2BERT reaches an accuracy of approximately 0.97, while HuBERT’s accuracy increases significantly to approximately 0.85. Importantly, the models’ generalization to other datasets remains unchanged, demonstrating the effectiveness and efficiency of few-shot fine-tuning. However, as the number of fine-tuning samples increases, HuBERT’s test accuracy on the WaveFake test set shows a declining trend, which is also observed for Wave2Vec2BERT.
It is important to note, however, that the efficiency and effectiveness of few-shot fine-tuning may vary across different datasets. As illustrated in Figures 2(c) and 2(d), which depict the fine-tuning results for Wave2Vec2BERT and HuBERT on Seed-TTS, the improvement in accuracy is less pronounced compared to the results on the OpenAI dataset. While the accuracy of both Wave2Vec2BERT and HuBERT does improve on Seed-TTS, the gains are not as significant as those observed for the OpenAI dataset. Additionally, the detection performance on other datasets decreases more noticeably when fine-tuning on Seed-TTS compared to OpenAI.
These findings suggest that the effectiveness of few-shot fine-tuning may depend on the specific characteristics of the dataset. Moreover, this also highlights its potential for tailored applications, such as personalized detection systems for a specific entity or individual, to enable more customized and practical applications.
Figure 2: Performance of few-shot fine-tuning for Wave2Vec2BERT and HuBERT with a varying number of few-shot audio samples from OpenAI and Seed-TTS, respectively. (a) Fine-tune Wave2Vec2BERT on OpenAI. (b) Fine-tune HuBERT on OpenAI; (c) Fine-tune Wave2Vec2BERT on Seed-TTS; and (d) Fine-tune HuBERT on Seed-TTS.
5 Discussion
AI-synthetized audio detection methods must be evaluated on diverse and advanced benchmarks. In our evaluation using the proposed dataset, most models perform well on standard TTS tools but suffer significant degradation when tested on the fake audios generated by the most advanced tool such as Voice Engine released by OpenAI. Therefore, we advocate for future research in audio deepfake detection to prioritize benchmarking against the latest and most advanced TTS technologies, which will lead to more robust and reliable detectors, as relying on high detection rates from outdated tools may create a false sense of generalization. Additionally, there is an urgent need to develop larger-scale training datasets comprising fake audio generated by cutting-edge TTS models to keep pace with rapid advancements in TTS technology and mitigate associated risks.
Limitations and future work. While our primary goal in proposing this dataset is to facilitate comprehensive evaluation, it remains relatively small in size and is primarily focused on English. A more in-depth analysis of detection performance across different languages and gender representations is crucial for a more comprehensive evaluation. These aspects are essential for future research to enhance the dataset’s applicability and generalizability. For future work, we also plan to: (1) incorporate additional AI-audio detection models, including those targeting advanced audio editing techniques designed to bypass detection systems; (2) explore innovative methods to further improve generalizability; and (3) address realistic challenges and risks in deploying the proposed method in real-world scenarios, such as evaluating the robustness of models against common or adversarial corruptions. These efforts will contribute to the development of more effective strategies to combat AI-generated audio threats.
Data license considerations.. Since our dataset is sourced from various models, each may be subject to different distribution licenses and usage restrictions. Throughout the data collection process, we ensured strict adherence to all relevant usage policies. We have also made the dataset accessible to everyone, either directly via the provided link or indirectly through the original sources. However, as these policies are frequently refined and updated, we will ensure that our published dataset remains in full compliance with the latest regulations. Additionally, we will reference the usage policies of the respective API providers to ensure that users are informed of any potential restrictions.
6 Conclusion
In this paper, we presented SONAR, a framework providing a comprehensive evaluation for distinguishing state-of-the-art AI-synthesized auditory content. SONAR introduces a novel evaluation dataset sourced from 9 diverse audio synthesis platforms, including leading TTS service providers and state-of-the-art TTS models. To the best of our knowledge, SONAR is the first platform that provides uniform, comprehensive, informative, and extensible evaluation of deepfake audio detection models. Leveraging SONAR, we conducted extensive experiments to analyze the generalizability limitations of current detection methods. We found that foundation models demonstrate stronger generalization capabilities, given their massive model size scale and pertaining data. We further explored the potential of few-shot fine-tuning to improve generalization and demonstrated its efficiency and effectiveness. We envision that SONAR will serve as a valuable benchmark to facilitate research in AI-audio detection and highlight directions for further improvement.
Acknowledgments and Disclosure of Funding
The authors thank the partial support from Fordham-IBM Research Fellowship and Fordham Faculty Research Grant.
Appendix A Appendix
A.1 Broad Impacts
Societal Risks. The rapid advancement of AI-generated content (AIGC) in audio and speech poses significant societal risks as it becomes more prevalent in audio and speech generation. As our work in benchmarking AI-synthesized audio detection demonstrates, the line between AI-generated audio and human speech is increasingly blurring, making it difficult for individuals to distinguish between synthetic and authentic voices. This raises serious concerns about spreading misinformation and fabricating narratives. AI-generated speeches could be used to impersonate public figures, spread false information, or even incite unrest by delivering provocative messages that appear authentic. For example, deepfake audios of political figures can be created to falsely represent their opinions or statements, potentially influencing public perception and affecting democratic processes.
Moreover, these technologies could be exploited to damage reputations or cause legal issues for individuals or organizations through fake endorsements or harmful statements. It is crucial for academia and industry to develop robust detection methods and ethical guidelines to prevent misuse of this technology and to educate the public about its capabilities and associated risks.
Positive Impacts. On the positive side, AI-synthesized audio/speech has the potential to revolutionize content creation in various sectors, including education, entertainment, and accessibility. In education, AI-synthesized audios and speeches enables production of customized content that meets diverse learning needs and languages, improving access and inclusivity. For entertainment, they can offer novel experiences by generating dynamic dialogues in games or virtual reality, enriching user engagement and creativity.
Furthermore, AI-synthesized audios and speeches also enhance accessibility by producing speech in various languages or dialects, bridging communication gaps and making information more accessible to non-native speakers or those with liabilities. Additionally, the technology can help preserve lesser-spoken languages and dialects at risk of extinction by creating archives of AI-generated speeches and narratives.
In conclusion, while AI-synthesized audios and speeches offer exciting opportunities for content creation and accessibility, it is essential to address the ethical and societal challenges associated with its use. Collaborative efforts among researchers, developers, and policymakers are crucial to leveraging AI-synthesized audio and speech benefits responsibly while mitigating its risks, ensuring the technology serves to enhance human communication and creativity positively and responsibly.
A.2 Implementation Details
Table 7: Hyperparemeters
config value
optimizer Adam
optimizer momentum $\beta_{1}=0.9$, $\beta_{2}=0.999$
weight decay 1e-4
epochs 40
warmup epochs 0
scheduler cosine decay
Table 7 presents the hyperparameters for training AASIST, RawNet2, RawGAT-ST, LCNN, and Spec.+ResNet. We train AASIST, RawNet2, and RawGAT-ST with a learning rate of 0.0001 and LCNN and Spec.+ResNet with a learning rate of 0.0003. The batch size for AASIST, RawNet2, RawGAT-ST, LCNN, and Spec.+ResNet are 64, 256, 32, 512, and 256, respectively. All input audios are resampled to a 16kHz sampling rate and converted into raw waveforms consisting of 64,000 samples (approximately 4 seconds). Audios longer than 4 seconds are randomly trimmed, while those shorter than 4 seconds are repeated and padded to meet the 4-second duration.
For the foundation models, two linear layers are added after the encoder’s output, with the hidden layer dimension matching the dimension of the encoder’s output. We fine-tune all foundation models on the Wavefake training dataset for 3 epochs using the Adam optimizer with a learning rate of 0.00001 and a weight decay of 0.0005.
For few-shot fine-tuning, models are fine-tuned for 30 epochs with a learning rate of 0.00001 and a weight decay of 0.00005.